HISTORY EVENTS IN AN INFLATIONARY BIG BANG COSMOLOGY

Time (s)  T_{CBR} (Kelvin) Reference  Event

THE QUANTUM GRAVITY ERA

1 × 10^{−43}  1 × 10^{32}  KT 72  Quantum limit of classical general relativity

THE INFLATION AND SYMMETRY BREAKING ERA

1 × 10^{−38}  1 × 10^{29}  KT 72  Limit of perturbative interaction thermalization of universe
1 × 10^{−35}  1 × 10^{28}  KT 72  Grand unification spontaneous symmetry breaking
1 × 10^{−34}  1 × 10^{27}  KT 274  Approximate start of inflation
1 × 10^{−32}  1 × 10^{27}  KT 274  Approximate start of reheating and end of inflation
1 × 10^{−11}  3 × 10^{15}  KT 72  Electroweak unification spontaneous symmetry breaking

THE QUARK-LEPTON ERA

2 × 10^{−7}  2 × 10^{13}  HA 353  Tauon anti-tauon annihilation
1 × 10^{−5}  2 × 10^{12}  KT 72  Formation of hadrons from quarks
7 × 10^{−5}  1 × 10^{12}  HA 353  Muon anti-muon annihilation
5 × 10^{−4}  4 × 10^{11}  KT 159, 281 By this time the universe has a baryon-antibaryon asymmetry

which results from post-inflationary B,C,CP violating processes

1 × 10^{−1} 3 × 10^{10} BS Neutral current weak interactions become too slow

and neutrinos decouple

1 × 10^{0} 1 × 10^{10} BS Charged current weak interactions become too slow

and the neutron to proton ratio freezes out

1 × 10^{1}  5 × 10^{9}  BS  Electron positron annihilation

THE RADIATION ERA

1 × 10^{2}  1 × 10^{9}  BS  Typical photon energies drop below the deuteron

binding energy and nucleosynthesis begins

1 × 10^{3} 4 × 10^{8} BS Particle energies drop below Coulomb barrier

energies and nucleosynthesis ends

4 × 10^{10}  6 × 10^{4}  KT 77  Matter density becomes equal to radiation density

THE MATTER ERA

4 × 10^{12}  3.5 × 10^{3}  KT 78  Electrons and protons recombine into hydrogen atoms
6 × 10^{12}  3.0 × 10^{3}  KT 80  Photon decoupling
2 × 10^{16}  20  HA 349  Start of the formation of galaxies
8 × 10^{16}  —  —  Approximate peak epoch of quasar activity
(X − 1.4 × 10^{17}) —  KT 12  Formation of the solar system
(X − 9.9 × 10^{16}) —  HA 390  Emergence of life on Earth
X ≈ 4.7 × 10^{17}  2.726  KT 12  Today

These values are rough estimates only. I have taken Ω_{0}=1, Λ_{0}=0, and h = 1.

BS=Boesgaard & Steigman, Annual Reviews of Astronomy and Astrophysics, 23, 322 HA=Harrison, Cosmology: The Science of the Universe KT=Kolb & Turner, The Early Universe